3. The amount of tea leaves in a can from a production line is normally distributed with μ = 110 grams and σ = 25 grams. A sample of 25 cans is to be selected. What is the probability that the sample mean will be between 100 and 120 grams?


1
Expert's answer
2022-03-10T18:20:38-0500

We have a normal distribution, "\mu=110, \sigma=25, n=25."

Let's convert it to the standard normal distribution, "z=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}" ;

"z_1=\cfrac{100-110}{25/\sqrt{25}}=-2, z_2=\cfrac{120-110}{25/\sqrt{25}}=2,"

"P(100<\bar{X}<120)=P(-2<Z<2)=P(Z<2)-P(Z<-2)="

"=0.97725-0.02275=0.9545" (from z-table).